# MAPGD: Multi-Agent Prompt Gradient Descent for Collaborative Prompt Optimization

### v2

@@
Abstract.
Prompt engineering is essential for fully leveraging large language models (LLMs), yet existing optimization methods often follow a single trajectory, leading to limited adaptability, gradient conflicts, and high computational overhead. We propose MAPGD (Multi-Agent Prompt Gradient Descent), a novel framework that reconceptualizes prompt optimization as a collaborative multi-agent process. MAPGD incorporates specialized agents focusing on orthogonal refinement dimensions—such as instruction clarity, example selection, format structuring, and stylistic adaptation—and coordinates their contributions through semantic gradient embedding, conflict detection, and fusion. To further enhance robustness and stability, MAPGD proposes two new mechanisms: Hypersphere-Constrained Gradient Clustering (HCGC), which enforces angular margin constraints to ensure intra-cluster compactness and inter-cluster separation, and Channel-Adaptive Agent Weighting (CAAW), which dynamically reweights agent contributions based on validation performance. Experiments on classification and reasoning benchmarks demonstrate that MAPGD consistently outperforms single-agent and random baselines in both accuracy and efficiency. Ablation studies validate the effectiveness of gradient fusion, agent specialization, and conflict resolution. Together, these contributions establish MAPGD as a unified, gradient-inspired, and interpretable framework for robust prompt optimization with theoretical convergence guarantees.
Key words and phrases:
Prompt Optimization, Multi-Agent Systems, Large Language Models, Gradient Descent
{CCSXML}
@@ Abstract.
1. Introduction
Figure 1. Overview of the Multi-Agent Prompt Gradient Descent for Collaborative Prompt Optimization.
@@ 1. Introduction
Overview of the Multi-Agent Prompt Gradient Descent for Collaborative Prompt Optimization.
Large Language Models (LLMs), trained on vast webscale corpora, have recently demonstrated remarkable generalization capabilities across diverse natural language processing (NLP) tasks Achiam et al. (2023); Bubeck et al. (2023). A central factor influencing their performance is the design of prompts, which serve as the primary interface for guiding model behavior. However, prompt construction remains largely a manual, trial-and-error process that demands substantial human labor Jiang et al. (2022) and domain expertise Reynolds and McDonell (2021); Zamfirescu-Pereira et al. (2023). This inefficiency highlights the urgent need for automatic or semi-automatic strategies to systematically generate or refine prompts. Such approaches can reduce manual burden, improve task performance, and provide interpretable optimization pathways.
Recent research has explored this problem through two primary directions. One line of work introduces auxiliary models or differentiable prompt embeddings to approximate gradient-based optimization Qin and Eisner (2021); Deng et al. (2022). Yet, these methods often assume access to the internal states or parameters of LLMs Shin et al. (2020); Lester et al. (2021), which is impractical in real-world API-only scenarios. Another line of work employs reinforcement learning or feedback-driven editing to discretely manipulate prompts Zhang et al. (2022); Zhou et al. (2022). While effective in some cases, these algorithms may produce incoherent prompts, incur high computational costs, or rely on unguided Monte Carlo search over the semantic space. Collectively, these limitations highlight the need for more robust, interpretable, and scalable frameworks for prompt optimization.
Addressing this gap, Pryzant et al. Pryzant et al. (2023) proposed ProTeGi, a nonparametric approach that mirrors gradient descent in textual space. ProTeGi generates natural language “gradients” from model errors and iteratively edits prompts in semantically opposite directions. To balance exploration and exploitation, it further integrates beam search and bandit-based candidate selection, achieving significant gains over earlier baselines. Nevertheless, ProTeGi operates under a single-agent paradigm, which restricts the diversity of optimization signals and leaves unresolved conflicts among competing refinement directions.
To overcome these limitations, we propose MAPGD (Multi-Agent Prompt Gradient Descent), a novel framework that reconceptualizes prompt optimization as a collaborative multi-agent process(Figure 1). MAPGD introduces a set of specialized agents, each dedicated to distinct refinement dimensions—such as instruction clarity, example selection, format structuring, and stylistic adaptation. To reconcile heterogeneous improvement signals, MAPGD employs a semantic gradient coordinator that embeds textual gradients into a shared vector space, enabling conflict detection, clustering, and fusion. In addition, we propose two novel mechanisms to further enhance robustness and stability: Hypersphere-Constrained Gradient Clustering (HCGC), which introduces angular margin constraints on the unit hypersphere to ensure compact intra-cluster coherence and inter-cluster separation, and Channel-Adaptive Agent Weighting (CAAW), which dynamically reweights agent contributions according to validation performance, amplifying effective optimization channels while suppressing noisy ones. Together, these innovations endow MAPGD with improved robustness, stability, and interpretability, enabling it to consistently outperform single-agent baselines such as ProTeGi across diverse optimization scenarios.
Beyond empirical improvements, MAPGD also provides a theoretical foundation for convergence. Under mild stochastic approximation assumptions, we prove that the multi-agent gradient coordination and fusion mechanisms preserve the convergence properties of classical gradient descent, yielding almost sure convergence to a local optimum at a sublinear rate of $\mathcal{O}(1/\sqrt{T})$, where $T$ denotes the number of iterations. This result establishes that, despite operating in a discrete, textual prompt space and coordinating multiple heterogeneous agents, MAPGD maintains stable and predictable optimization dynamics, bridging practical prompt refinement with rigorous theoretical guarantees.
Figure 2. Workflow of MAPGD. Starting from an initial prompt $p$, specialized agents propose diverse pseudo-gradients $\{g,g^{\prime},g^{\prime\prime},g^{\prime\prime\prime},\ldots\}$ (top). These gradients are semantically embedded and projected onto the unit hypersphere, then clustered with angular-margin constraints to form coherent groups (middle). Clustered directions are fused via LLM-driven synthesis into composite gradients $G$, which generate refined candidate prompts $\{P_{1},P_{2},\ldots,P_{n}\}$ (bottom). The best candidates are iteratively selected and fed back to continue the prompt optimization loop.
@@
Workflow of MAPGD. Starting from an initial prompt $p$, specialized agents propose diverse pseudo-gradients $\{g,g^{\prime},g^{\prime\prime},g^{\prime\prime\prime}$…\} (top). These gradients are semantically embedded and projected onto the unit hypersphere, then clustered with angular-margin constraints to form coherent groups (middle). Clustered directions are fused via LLM-driven synthesis into composite gradients $G$, which generate refined candidate prompts $\{P_{1},P_{2},\ldots,P_{n}\}$ (bottom). The best candidates are iteratively selected and fed back to continue the prompt optimization loop.
2. Related Work
Prompt Learning and Optimization.
Early efforts relied on manual engineering, which lacks scalability. Automated methods include reinforcement learning Deng et al. (2022), evolutionary search Fernando et al. (2023), and Bayesian optimization Sahoo et al. (2025). Continuous prompt tuning (e.g., prefix tuning Li and Liang (2021), soft prompts Lester et al. (2021)) optimizes embeddings but sacrifices interpretability. MAPGD instead pursues interpretable natural language optimization with structured feedback.
Multi-Agent Collaboration.
Multi-agent systems are well-studied in RL, distributed AI, and game theory. In NLP, multi-agent debate Liang et al. (2023); Du et al. (2023), cooperative reasoning, and collaborative generation Hong et al. (2024) show that role specialization outperforms monolithic optimization. MAPGD applies this principle by assigning agents to distinct prompt dimensions and coordinating them via gradient fusion.
Gradient-Inspired Optimization.
ProTeGi Pryzant et al. (2023) and similar approaches approximate prompt gradients through self-feedback. While effective, single-agent descent struggles with limited signal diversity and inherent conflict resolution challenges Xiang et al. (2025). A monolithic agent can generate self-contradictory pseudo-gradients, as different refinement dimensions are not explicitly disentangled. For instance, to correct a misclassification, a single agent might propose a gradient to add more illustrative examples, while an implicit counter-gradient aims to make the instructions more concise. This creates a direct conflict in prompt length and structure, potentially leading to suboptimal or oscillating optimization trajectories. MAPGD extends this paradigm by embedding gradients semantically, explicitly detecting such conflicts, and fusing them coherently through LLM-driven synthesis.
Geometric and Adaptive Mechanisms.
Contrastive learning and angular margin constraints are effective in representation learning. Wang and Isola formalize how contrastive losses enforce alignment and uniformity, providing a theoretical lens Wang and Isola (2020). This principle of hyperspherical separation to enhance discriminability spans domains. In face recognition, margin-based softmax losses such as SphereFace Liu et al. (2017), CosFace Wang et al. (2018), and ArcFace Deng et al. (2019) enforce angular margins between identities. Chen et al. Chen et al. (2021) introduce a distance polarization regularizer for large-margin contrastive learning, pushing inter-cluster samples apart while keeping intra-cluster representations tight. Inspired by these advances, our Hypersphere-Constrained Gradient Clustering (HCGC) enforces intra-cluster compactness and inter-cluster separation for semantic gradients, complemented by Channel-Adaptive Agent Weighting (CAAW) adjusting agent contributions. Together, these modules enhance multi-agent gradient descent with geometry-aware clustering and adaptive coordination.
In summary, MAPGD unifies prompt learning, multi-agent collaboration, gradient descent, and geometry-inspired optimization into a scalable and interpretable framework.
3. Methodology
3.1. Framework Overview
MAPGD formulates prompt optimization as a hybrid discrete–continuous gradient descent process in the natural language space. Unlike continuous embedding–based methods such as prefix or soft prompt tuning Li and Liang (2021); Lester et al. (2021), MAPGD explicitly manipulates interpretable textual prompts while leveraging gradient-inspired feedback for iterative refinement (Figure 2).
At each iteration, specialized agents analyze misclassified or suboptimal examples and produce pseudo-gradients in natural language. These gradients are semantically embedded, clustered, and adaptively fused to mitigate conflicts and ensure coherent optimization. The resulting fused gradients are applied to generate successor prompts, which are then evaluated and filtered under limited budgets. Iterative repetition yields a dynamically evolving prompt trajectory that balances exploration of diverse directions with exploitation of promising refinements.
Formally, the optimization objective is defined as:
where $\ell(\cdot,\cdot)$ denotes a task-specific loss function. The goal is:
$p^{*}=\arg\min_{p}F(p).$ (2)
Unlike stochastic gradient descent (SGD), where gradients are continuous, MAPGD constructs textual pseudo-gradients:
$\nabla F(p^{(t)})\approx g^{(t)},$ (3)
which act as semantic analogues of numerical gradients.
3.2. Specialized Prompt Agents
Each agent specializes in a distinct dimension of prompt optimization, analogous to orthogonal gradient directions. For example: $A_{1}$: instruction clarity ($g_{1}$), $A_{2}$: example selection ($g_{2}$), $A_{3}$: format specification ($g_{3}$), $A_{4}$: stylistic refinement ($g_{4}$), and optionally a generic agent for broad improvements. At iteration $t$, the gradient set is:
$G^{(t)}=\{g_{1}^{(t)},g_{2}^{(t)},\dots,g_{K}^{(t)}\}.$
This decomposition supports parallel exploration of diverse refinements, alleviating local minima issues inherent in single-agent approaches.
3.3. Hypersphere-Constrained Gradient Clustering
Inspired by hyperspherical contrastive learning Wang et al. (2018); Liu et al. (2017), our HCGC mechanism is designed to enhance the semantic consistency of multi-agent gradients by enforcing intra-cluster compactness and inter-cluster separation in a normalized hypersphere space. This approach addresses the challenge of conflicting pseudo-gradients generated by heterogeneous agents, ensuring that semantically coherent optimization directions are emphasized while contradictory ones are disentangled. The overall process and the geometric intuition of our method are illustrated in Figure 3.
Gradient Embedding on Hypersphere.
Given a set of agent-generated gradients $G^{(t)}=\{g_{1}^{(t)},g_{2}^{(t)},\dots,g_{K}^{(t)}\}$, each gradient is encoded into a d-dimensional vector $v_{k}^{(t)}=\phi(g_{k}^{(t)})$, where $\phi$ denotes a pre-trained language encoder. For this purpose, we utilize a Sentence Transformer model, such as all-MiniLM-L6-v2, which provides compact and semantically informative embeddings. To eliminate scale variance, we normalize each embedding onto the unit hypersphere:
The similarity between two gradients is then computed as the cosine similarity. Gradients are considered to be in conflict if their angular distance exceeds a predefined threshold, $\theta_{conflict}$:
where $\Delta(\cdot,\cdot)$ denotes the angular distance on the hypersphere.
Figure 4 illustrates a concrete example in the sentiment classification task, where different agents generate multiple pseudo-gradients. While some gradients form semantically coherent clusters (intra-cluster compactness), others introduce conflicts that must be separated through angular margin constraints.
Angular Margin Constraint.
To mitigate conflicts and ensure clear separation between semantic clusters, HCGC enforces an angular margin constraint on the unit hypersphere. Let $\hat{v}_{k}$ denote a normalized gradient vector, $u_{i}$ the centroid of its assigned cluster $i$, and $u_{j}$ the centroid of any other cluster $j$. We define the corresponding angles as
$\alpha=\Delta(\hat{v}_{k},u_{i}),\quad\beta=\Delta(\hat{v}_{k},u_{j}).$
A standard clustering assignment only requires $\alpha<\beta$. However, to enforce stronger separation, we introduce a margin scaling factor $n\geq 1$:
$n\cdot\alpha<\beta.$
Since $\cos(\theta)$ is monotonically decreasing for $\theta\in[0,\pi]$, this condition is equivalently expressed in terms of cosine similarity:
$\cos(n\cdot\alpha)>\cos(\beta),\quad n\geq 1.$ (6)
This constraint ensures that $\hat{v}_{k}$ is not only closer to its own centroid $u_{i}$, but also significantly more aligned with it than with any other centroid $u_{j}$. Geometrically, as illustrated in Figure 3 (c), this confines $\hat{v}_{k}$ to a scaled angular neighborhood around $u_{i}$, effectively creating a “cone of acceptance” narrower than the standard Voronoi region on the hypersphere. By enforcing this margin, HCGC simultaneously enhances inter-cluster separation and intra-cluster compactness, yielding more stable and reliable gradient fusion. During clustering, any gradient $\hat{v}_{k}$ violating this condition with respect to some $u_{j}$ is considered for reassignment to a cluster where the margin holds; otherwise, it remains in its original cluster.
Figure 3. Details of Hypersphere-Constrained Gradient Clustering

### v3

@@
Abstract
Prompt engineering is crucial for fully leveraging large language models, yet existing prompt optimization methods often rely on a single refinement trajectory, leading to instability, conflicting update signals, and inefficient use of query budgets. We propose Multi-Agent Prompt Gradient Descent (MAPGD), a gradient-inspired framework that coordinates multiple complementary prompt editing signals for robust and interpretable optimization. MAPGD decomposes prompt refinement into orthogonal dimensions (e.g., instruction clarity, example selection, format, and style) and aggregates textual pseudo-gradients via semantic embedding, conflict-aware clustering, and adaptive fusion. To improve robustness, we introduce HCGC, which enforces angular separation between conflicting directions, and CAAW, which dynamically calibrates editing contributions based on validation feedback. Experiments on diverse classification and reasoning benchmarks show that MAPGD achieves consistent gains in accuracy over strong baselines, while ablation results indicate that coordinated gradient fusion and adaptive weighting are critical to stable optimization. Together, these findings suggest that MAPGD offers a robust and interpretable framework for automatic prompt optimization.
MAPGD: Multi-Agent Prompt Gradient Descent for Collaborative Prompt Optimization
@@ Abstract
1 Introduction
Large Language Models (LLMs), trained on vast web-scale corpora, have demonstrated remarkable generalization capabilities across a wide range of natural language processing (NLP) tasks, including classification, reasoning, and generation Achiam et al. (2023); Bubeck et al. (2023). A critical factor underlying this performance is the design of prompts, which serve as the primary interface for steering model behavior. Despite their importance, prompts are still predominantly crafted through manual trial-and-error, requiring substantial human effort Jiang et al. (2022) and domain expertise Reynolds and McDonell (2021); Zamfirescu-Pereira et al. (2023). This reliance on manual design limits scalability and motivates the development of automatic or semi-automatic prompt optimization methods that are both effective and interpretable.
Existing approaches to prompt optimization largely fall into two categories. The first line of work approximates gradient-based optimization through auxiliary models or differentiable prompt representations Qin and Eisner (2021); Deng et al. (2022). While theoretically appealing, these methods often assume access to internal model states or parameters Shin et al. (2020); Lester et al. (2021), making them impractical in real-world, API-only settings. The second line of work formulates prompt optimization as a discrete search or reinforcement learning problem, using feedback-driven edits or Monte Carlo exploration Zhang et al. (2022); Zhou et al. (2022). Although effective in certain cases, these methods can suffer from high computational cost, unstable optimization trajectories, and limited interpretability. Collectively, these limitations highlight the need for more robust, scalable, and transparent prompt optimization frameworks.
To bridge this gap, Pryzant et al. Pryzant et al. (2023) proposed ProTeGi, a non-parametric framework that reinterprets prompt optimization as gradient descent in natural language space. ProTeGi generates textual pseudo-gradients from model errors and iteratively refines prompts using beam search and bandit-based selection, achieving strong empirical performance. However, ProTeGi relies on a single-agent optimization framework that aggregates heterogeneous refinement signals into a unified trajectory. Such entanglement may give rise to implicit objective conflicts, thereby impairing optimization stability and convergence quality.
A natural extension is to decompose prompt optimization into multiple specialized perspectives. Multi-agent formulations enable parallel exploration of heterogeneous refinement dimensions, increasing signal diversity and reducing premature convergence. However, introducing multiple agents also makes semantic conflicts inevitable: different agents may propose mutually incompatible updates, such as simultaneously expanding illustrative examples while enforcing concise instructions. Without explicit coordination mechanisms, such conflicts can destabilize optimization and undermine the benefits of multi-agent collaboration.
In this work, we propose MAPGD, a unified framework that addresses these challenges by integrating multi-agent specialization with principled coordination mechanisms (Figure 1). As illustrated in Figure 1(a), MAPGD decomposes prompt optimization into multiple orthogonal refinement dimensions and assigns each to a specialized agent. Each agent independently analyzes model errors and produces textual pseudo-gradients that capture targeted improvement directions.
However, pseudo-gradients produced by different agents may encode semantically conflicting refinement intents. To explicitly disentangle such conflicts, MAPGD proposes Hypersphere-Constrained Gradient Clustering (HCGC), which embeds agent-generated gradients into a shared semantic space and separates incompatible optimization directions via angularly constrained clustering. This design promotes semantic consistency within clusters while preventing destructive interference across conflicting refinement signals.
While HCGC addresses conflicts at the directional level, agent contributions within the same semantic cluster may still vary in reliability. MAPGD therefore incorporates CAAW, which dynamically reweights agent channels based on validation performance. By emphasizing consistently effective signals and attenuating noisy or unstable ones, CAAW enables robust aggregation of non-conflicting gradients, leading to stable and effective prompt refinement.
In summary, our contributions are three-fold.
(1) We propose MAPGD, a multi-agent pseudo-gradient framework for stable and interpretable prompt optimization.
(2) We introduce HCGC and CAAW to disentangle conflicting refinement signals and adaptively aggregate reliable agent contributions.
(3) We establish theoretical convergence guarantees, showing that MAPGD preserves classical gradient descent behavior with almost sure convergence to a local optimum at rate $\mathcal{O}(1/\sqrt{T})$.
Figure 1: Overview of MAPGD. MAPGD performs prompt optimization through collaborative multi-agent pseudo-gradients. Agent-generated gradients are embedded and clustered via HCGC to resolve semantic conflicts, and adaptively weighted by CAAW according to validation performance. The resulting fused gradients iteratively refine the prompt through candidate generation and selection.
2 Related Work
Prompt Learning and Optimization.
Early prompt engineering relied on manual design, which is labor-intensive and difficult to scale. Automated approaches have since been explored, including reinforcement learning Deng et al. (2022), evolutionary search Fernando et al. (2023), and Bayesian optimization Sahoo et al. (2025). Continuous prompt tuning methods, such as prefix tuning Li and Liang (2021) and soft prompts Lester et al. (2021), optimize prompts in embedding space but often sacrifice interpretability and require access to model internals. MAPGD operates in natural language space, preserving interpretability while enabling feedback-driven optimization.
Gradient-Inspired Prompt Optimization.
Recent work has framed prompt optimization as a gradient-like process in discrete language space. ProTeGi Pryzant et al. (2023) generates natural language pseudo-gradients from model errors and iteratively refines prompts via beam search and bandit-based selection. However, as a single-agent method, ProTeGi entangles heterogeneous refinement objectives and lacks explicit mechanisms for resolving conflicting optimization signals, which can lead to unstable or suboptimal convergence. This limitation motivates approaches that can disentangle and coordinate diverse refinement directions.
Multi-Agent Collaboration.
Multi-agent formulations enable parallel exploration of diverse perspectives and have been widely studied in reinforcement learning and distributed AI. In NLP, multi-agent debate Liang et al. (2023); Du et al. (2023) and collaborative generation Hong et al. (2024) demonstrate the benefits of role specialization. However, increased diversity makes semantic conflicts inevitable, as different agents may propose incompatible updates. MAPGD adopts a multi-agent design with specialized roles, while explicitly addressing the resulting coordination and conflict resolution challenges.
Geometric and Adaptive Coordination.
Resolving semantic conflicts requires structured inductive biases beyond heuristic fusion. In representation learning, contrastive objectives and angular margin constraints enforce intra-cluster compactness and inter-cluster separation Wang and Isola (2020); Liu et al. (2017); Wang et al. (2018); Deng et al. (2019). Chen et al. Chen et al. (2021) further enhance large-margin contrastive learning via distance polarization regularization. Inspired by these advances, MAPGD introduces HCGC for geometry-aware gradient coordination, complemented by CAAW to adaptively calibrate agent contributions.
Concurrent Work.
Concurrent studies such as MAPRO Zhang et al. (2025) and MA-SAPO Seo et al. (2025) also explore multi-agent prompt optimization under different assumptions, including explicit interaction structures or external memory components. These settings are not directly aligned with our budgeted, task-level optimization protocol. A detailed comparison is provided in Appendix E.
In summary, MAPGD unifies prompt optimization, gradient-inspired refinement, multi-agent collaboration, and geometry-aware coordination into an interpretable framework that explicitly addresses semantic conflicts.
3 Methodology
3.1 Framework Overview
MAPGD formulates prompt optimization as a hybrid discrete–continuous gradient descent process operating directly in natural language space. Unlike continuous embedding-based approaches such as prefix tuning or soft prompt optimization Li and Liang (2021); Lester et al. (2021), MAPGD preserves interpretability by explicitly manipulating textual prompts while still leveraging gradient-inspired feedback signals for iterative refinement. As illustrated in Figure 1, MAPGD frames prompt optimization as a collaborative multi-agent process, in which specialized agents focus on orthogonal refinement dimensions and produce textual pseudo-gradients that guide prompt updates.
Building on this formulation, MAPGD instantiates the optimization procedure through an iterative coordination loop. At each iteration, specialized agents independently examine the current prompt and available task feedback, and generate natural language pseudo-gradients that reflect distinct refinement perspectives. These gradients are embedded into a shared semantic space, where potential conflicts are explicitly identified and mitigated via geometry-aware clustering and adaptive fusion. The resulting fused gradients guide the generation of successor prompts, which are subsequently evaluated and filtered under a budgeted selection mechanism. Repeating this process yields an evolving optimization trajectory that systematically balances exploration across heterogeneous refinement directions with exploitation of effective updates.
Formally, the prompt optimization objective is defined as
where $\ell(\cdot,\cdot)$ denotes a task-specific loss function and $M(x;p)$ represents the model output conditioned on prompt $p$. The optimization goal is
$p^{*}=\arg\min_{p}F(p).$ (2)
Since gradients in this discrete textual space are not directly accessible, MAPGD constructs natural language pseudo-gradients,
$\nabla F(p^{(t)})\approx g^{(t)},$ (3)
which serve as semantic analogues of numerical gradients in classical optimization.
3.2 Specialized Prompt Agents
MAPGD employs multiple specialized agents, each responsible for a distinct and approximately orthogonal dimension of prompt refinement. Typical agent roles include instruction clarity, example selection, output formatting, and stylistic refinement, optionally complemented by a generic agent for broad exploratory updates. At iteration $t$, the resulting pseudo-gradient set is
$G^{(t)}=\{g_{1}^{(t)},g_{2}^{(t)},\dots,g_{K}^{(t)}\}.$
By decomposing prompt optimization into specialized perspectives, MAPGD enables parallel exploration of heterogeneous refinement directions, alleviating the limited signal diversity and premature convergence commonly observed in single-agent optimization.
3.3 Hypersphere-Constrained Gradient Clustering
Inspired by hyperspherical metric learning Wang et al. (2018); Liu et al. (2017), HCGC resolves semantic conflicts among heterogeneous pseudo-gradients by enforcing geometry-aware separation on a normalized hypersphere. Instead of assuming all agent proposals are mutually compatible, HCGC explicitly detects semantic disagreement and clusters gradients into coherent refinement directions before fusion.
Gradient Embedding and Conflict Signal.
Given agent-generated pseudo-gradients $G^{(t)}$, each $g_{k}^{(t)}$ is encoded into a $d$-dimensional semantic vector $v_{k}^{(t)}=\phi(g_{k}^{(t)})$ and normalized onto the unit hypersphere:
Semantic similarity is measured by cosine similarity,
where $\Delta(\cdot,\cdot)$ denotes the angular distance on the hypersphere. To make “conflict” explicit and measurable, we define two gradients as being in semantic conflict if their similarity falls below a threshold:
$\mathrm{conflict}(\hat{v}_{i},\hat{v}_{j})\;\Leftrightarrow\;\hat{v}_{i}^{\top}\hat{v}_{j}<\theta_{\text{conflict}}.$ (6)
Intuitively, low cosine similarity indicates substantially different (and potentially incompatible) refinement intents, motivating explicit separation in clustering.
Clustering and Adaptive Cluster Cardinality.
We perform cosine K-means on the hypersphere to group semantically coherent directions. Rather than fixing the number of clusters, we adopt a bounded adaptive strategy:
$K^{(t)}=\min\big(|G^{(t)}|,\;K_{\max}\big),$ (7)
where $|G^{(t)}|$ is the number of gradients at iteration $t$ and $K_{\max}$ is an upper bound. This prevents over-fragmentation when gradients are few, while allowing multiple distinct directions when agent proposals are diverse.
Angular Margin Constraint.
Initial clustering may still yield weakly separated clusters when conflicting gradients coexist. HCGC therefore refines assignments with an angular margin constraint. Let $u_{i}$ be the centroid of $\hat{v}_{k}$’s assigned cluster and $u_{j}$ any other centroid. Define $\alpha=\Delta(\hat{v}_{k},u_{i})$ and $\beta=\Delta(\hat{v}_{k},u_{j})$. While standard assignment requires $\alpha<\beta$, HCGC strengthens it by a margin scale $n\geq 1$:
$n\cdot\alpha<\beta,$ (8)
equivalently,
$\cos(n\cdot\alpha)>\cos(\beta).$ (9)
Gradients that violate the margin are reassigned to clusters where the constraint holds, yielding more compact and better-separated semantic groups.
Fusion of Clustered Gradients.
Finally, gradients within each cluster are fused into a single representative refinement direction via LLM-guided synthesis. To avoid redundant updates, we apply a diversity filter to discard fused gradients overly similar to those already selected, controlled by $\theta_{\text{diversity}}$.
Figure 2: Effect of hypersphere-constrained angular margin in HCGC. (a) Clustering without angular margin constraints may entangle semantically conflicting pseudo-gradients, leading to ambiguous cluster boundaries. (b) Enforcing an angular margin on the unit hypersphere improves intra-cluster compactness and inter-cluster separation, effectively isolating conflicting refinement directions.
3.4 Channel-Adaptive Agent Weighting
Although HCGC resolves structural conflicts between semantic gradient directions, gradients within the same cluster may still vary in reliability due to heterogeneous agent expertise and historical effectiveness. To account for this variability, we introduce CAAW, a reliability-aware coordination mechanism that adaptively calibrates agent contributions during gradient fusion. CAAW is inspired by channel-wise recalibration mechanisms in neural representations Hu et al. (2018), but operates at the level of agent-generated textual gradients.
Adaptive Weight Estimation.
Let $s_{k}^{(t)}$ denote the validation performance gain attributed to gradient $g_{k}^{(t)}$ at iteration $t$, estimated from historical agent performance statistics. CAAW converts these scalar signals into normalized reliability weights via a softmax transformation:
where $C$ denotes the set of gradients within the same semantic cluster and $\lambda$ controls the sharpness of the weighting distribution. This formulation assigns higher weights to agents that have consistently produced effective refinements, while down-weighting less reliable channels.
Weight-Conditioned LLM Fusion.
Rather than explicitly performing an arithmetic weighted sum over textual gradients, CAAW injects the estimated weights into the LLM-driven fusion operator as conditioning signals. Concretely, the fusion operator
$g_{\text{fused}}^{(t)}=\Psi\big(\{g_{k}^{(t)}\}_{k\in C}\,\big|\,\{w_{k}^{(t)}\}_{k\in C}\big)$ (11)
receives both the set of clustered gradients and their corresponding reliability weights, which are encoded in the fusion prompt to bias the synthesis process toward more trustworthy agent suggestions. In this way, agent weights influence the fused gradient implicitly through prompt-level emphasis, rather than through explicit numerical combination.
Stability Implications.
By conditioning gradient fusion on adaptive reliability weights, CAAW suppresses noisy or redundant updates while amplifying consistently effective refinement channels. This weight-aware fusion mechanism reduces variance in the optimization trajectory and complements HCGC by addressing intra-cluster reliability differences, thereby promoting stable convergence toward semantically robust prompt candidates.
3.5 Candidate Generation and Selection
Fused gradients generate successor prompts $\{p^{\prime}_{1},\dots,p^{\prime}_{n}\}$. MAPGD applies lightweight filtering and optional bandit-based evaluation Audibert and Bubeck (2010) to keep the computation budgeted. Unlike prior work that depends heavily on bandit selection, MAPGD attains robustness and interpretability mainly via conflict-aware clustering and adaptive weighting across agents. Overall algorithm can be found in Algorithm 1.
Algorithm 1 MAPGD with HCGC and CAAW
3:for $t=1$ to $R$ do
4:  $M_{t}\leftarrow\textsc{SampleMiniBatch}(D_{\text{train}},b)$
5:  $\mathcal{G}_{t}\leftarrow\textsc{GeneAgentGrads}(\{A_{i}\},M_{t})$ $\triangleright$ Alg. 2
7:  $F_{t}\leftarrow\textsc{CAAW}(\tilde{\mathcal{G}}_{t},D_{\text{dev}})$ $\triangleright$ Alg. 4
@@ 3.5 Candidate Generation and Selection
12:  end if
13:end for
14:return Best prompt over $\{p_{\star}^{(1)},\dots,p_{\star}^{(t)}\}$
3.6 Theoretical Convergence
We provide a theoretical analysis of MAPGD in Appendix G. Here, we summarize the key assumptions and main implications to clarify how the proposed framework relates to classical stochastic optimization.
Specifically, our analysis relies on three standard conditions: (i) bounded second moments of agent-generated pseudo-gradients, (ii) smoothness (or Lipschitz continuity) of the empirical loss function in the semantic embedding space, (iii) unbiased sampling induced by the bandit-based candidate selection mechanism. Under these assumptions, MAPGD admits an interpretation as a stochastic approximation process and achieves a sublinear convergence rate of $\mathcal{O}(1/\sqrt{T})$ in both convex and non-convex settings.
Intuitively, the core components of MAPGD play complementary roles in ensuring stable optimization. HCGC aggregates semantically aligned refinement directions and mitigates conflicting updates, reducing variance in the fused pseudo-gradient. CAAW further emphasizes agents that have demonstrated consistent utility across iterations. Finally, bandit-based candidate selection maintains sufficient exploration while preventing systematic bias. Together, these mechanisms ensure that the overall optimization dynamics remain stable and progressively improve prompt quality over iterations.
4 Experiments
Experiments evaluate MAPGD in terms of effectiveness, efficiency, and robustness. For fair comparison, we reproduce ProTeGi under identical settings and further conduct ablations and token consumption analysis. A real-world text-generation case study is also included (Appendix I) to demonstrate applicability beyond benchmarks.
4.1 Experimental Details
Datasets. Following ProTeGi’s protocol, we evaluate on four classification benchmarks (Jailbreak, Ethos, LIAR, Sarcasm) and three arithmetic reasoning datasets (GSM8K, AQUARAT, SVAMP). For each dataset, we sample 50/150 instances for dev/test and report test F1 (classification) or accuracy (reasoning). Dataset details are in Appendix A.
Setup. We use the January 2023 gpt-3.5-turbo via Azure OpenAI unless stated otherwise (temperature $0.0$ for classification). Results are averaged over three runs; F1 is computed by max-pooling over the final prompt beam. All methods share identical initial prompts, data, and random seeds, with a fixed budget of T=10T{=}10 and default hyperparameters. Unless stated otherwise, MAPGD follows Section 3.3 with $\phi{=}\texttt{all-MiniLM-L6-v2}$, $\theta_{\text{conflict}}{=}0.3$, clustering threshold $0.7$, $\theta_{\text{diversity}}{=}0.7$, and UCB selection with 80 evaluations; $\lambda{=}1$ (Appendix D). All experiments in Section 4.2 use these modules unless ablated.
Baselines. We primarily compare against ProTeGi Pryzant et al. (2023) under identical query budgets and initialization, and additionally include MC search, RL-style edit methods, AutoGPT, and three reasoning baselines (InsZero, Instinct, PromptWizard). Full descriptions are in Appendix B.
4.2 Experimental Results
Figure 3: Test performance (F1 score) vs. API query budget per prompt candidate across four benchmark tasks (evaluated up to 50 queries for fair comparison with ProTeGi).
Classification benchmarks. Figure 3 reports test F1 under different per-candidate evaluation budgets (up to 50 queries, following ProTeGi for fairness). MAPGD consistently achieves the best performance across all four datasets, with most gains emerging within the first 30–40 evaluations. The improvements are most pronounced on conflict-prone tasks, suggesting that explicit conflict disentanglement and adaptive coordination are critical when refinement signals disagree: on Jailbreak, MAPGD reaches 0.880 at 50 queries (vs. 0.833 for ProTeGi), and on LIAR it attains 0.710 (vs. 0.660). We attribute these gains to HCGC, which clusters compatible pseudo-gradients before fusion to reduce semantic interference, and CAAW, which down-weights unreliable agents to prevent noisy edits from dominating early iterations, thereby improving budget efficiency. On the more saturated Ethos dataset, MAPGD still improves to 0.980 (vs. 0.965), indicating that coordinated multi-agent feedback can extract residual refinements even near a performance ceiling. We follow the 50-query budget for strict comparability; extended-budget results are deferred to Appendix C.
Arithmetic reasoning benchmarks. Table 1 evaluates MAPGD on GSM8k, AQUARAT, and SVAMP, compared against InsZero, Instinct, and PromptWizard. MAPGD achieves consistent gains across all three datasets, most notably on GSM8k (93.5%), improving over the strongest baseline by +3.5 points. We hypothesize that the benefit on reasoning stems from the same coordination principle: specialized agents propose complementary edits (e.g., step-by-step constraints, format strictness, and exemplar guidance), while HCGC/CAAW suppress incompatible or low-yield directions, leading to more stable chain-of-thought prompting. MAPGD also improves the best baseline on AQUARAT (60.3% vs. 58.2%) and SVAMP (84.1% vs. 82.3%), confirming transfer beyond classification.
Method GSM8k AQUARAT SVAMP
InsZero 74.2 54.3 79.5
Instinct 74.5 54.7 81.0
PromptWizard 90.0 58.2 82.3